The total revenue function for two goods is given by the equation TR = 36 x-3x²+ 56y- 4y².find the number of unit of each goods which must be sold. if profit is to be maximize when the firm subject to a budget constraint 5 x + 10 y = 80
1
Expert's answer
2021-10-27T13:55:40-0400

"TR=36x+3x\u00b2+56y-4y\u00b2"

Budget constant "5y+10y=80"

"x+2y=16\\x=16-2y"

"TR=36(16-2y)+3(16-2y)\u00b2+56y-4y\u00b2"

"\frac{∆TR}{∆Y}=-72y+6(16-2y)+56+8y"

"\frac{∆TR}{∆Y}=-104y+248=0"

"y=\frac{248}{104}"

"x=16-\frac{496}{104}=\frac{1168}{104}"

"y=2.385\\x=11.231"